GONALITY OF MODULAR CURVES IN CHARACTERISTIC p

BJORN POONEN

Abstract. Let k be an algebraically closed field of characteristic p. Let X(p^{e}; N) be the curve parameterizing elliptic curves with full level N structure (where p ∤ N) and full level p^{e} Igusa structure. By modular curve, we mean a quotient of any X(p^{e}; N) by any subgroup of ((ℤ/p^{e}ℤ)^{×} × SL_{2}(ℤ/Nℤ)) /{±1}. We prove that in any sequence of distinct modular curves over k, the k-gonality tends to infinity. This extends earlier work, in which the result was proved for particular sequences of modular curves, such as X_{0}(N) for p ∤ N. As an application, we prove the function field analogue of a uniform boundedness conjecture for the image of Galois on torsion of elliptic curves.

1. Introduction

1.1. Gonality. The gonality γ_{k}(X) of a curve^{1} X over a field k is the smallest possible degree of a dominant rational map X −−→ ℙ_{k}^{1}. For any field extension L of k, we define also the L-gonality γ_{L}(X) of X as the gonality of X_{L} := X ×_{k} L. General facts about gonality (mostly well-known) are gathered in Proposition A.1 in the appendix.

1.2. Modular curves. Our goal is to obtain lower bounds on the gonality of modular curves. By Proposition A.1(ii), it suffices to consider k = ℂ and k = \overline{𝔽}_{p} (an algebraic closure of 𝔽_{p}) for each prime p.

Suppose N is a positive integer not divisible by the characteristic of k. Choose a primitive N-th root of unity ζ. Let X(N) be the smooth projective model of the (possibly coarse) moduli space parameterizing triples (E, P, Q) where P, Q ∈ E are a basis for E[N] with Weil pairing e_{N}(P, Q) = ζ. For k = ℂ, we can describe X(N)(ℂ) alternatively as the quotient of an extended upper half plane by a finite-index subgroup of PSL_{2}(ℤ). More generally, any congruence subgroup G ≤ PSL_{2}(ℤ) gives rise to a curve X_{G} over ℂ. Abramovich proved:

Theorem 1.1 ([Abr96]). Let D = (PSL_{2}(ℤ) : G). Then γ_{ℂ}(X_{G}) ≥ \frac{7}{800}D.

Remark 1.2. As mentioned in [Abr96], combining Theorem 1.1 with the genus bound g −1 ≤ D/12 [Shi94, Proposition 1.40] yields

γ_{ℂ}(X_{G}) ≥ \frac{21}{200}(g − 1).

Date: April 16, 2006.

2000 Mathematics Subject Classification. Primary 14G35; Secondary 14H51, 11G18.

Key words and phrases. gonality, modular curve, Igusa curve, image of Galois.

This research was supported by NSF grant DMS-0301280 and the Miller Institute for Basic Research in Science.

1All our curves are assumed to be geometrically integral.

The proof of Theorem 1.1 makes use of a lower bound on the leading nontrivial eigenvalue of the noneuclidean Laplacian; this bound has been improved since 1996, so the constants 7/800 and 21/200 can be improved too. See also [BGJGP05, §4.3] for some further results.

In characteristic p, one has other kinds of modular curves, involving level structure where p divides the level. If q = p^{e} for some e ∈ ℤ_{≥1}, the Igusa curve of level q is the smooth projective model Ig(q) of the curve over 𝔽_{p} parameterizing pairs (E, R) where E is an ordinary elliptic curve, and R is a generator of the kernel of the degree-q Verschiebung isogeny V_{q}: E^{(q)} → E, where E^{(q)} is the elliptic curve obtained by raising all the coefficients of a model of E to the q-th power. Given N not divisible by p, and e ∈ ℤ_{≥1}, we can define also a hybrid modular curve X(p^{e}; N) over \overline{𝔽}_{p} parameterizing (E, R, P, Q), with R generating ker V_{q} and P, Q ∈ E[N] as above.

The group G_{p^{e}N} := (ℤ/p^{e}ℤ)^{×} × SL_{2}(ℤ/Nℤ) acts on X(p^{e}; N). The kernel of the action is {±1} embedded diagonally in G_{p^{e}N}. For any subgroup G ≤ G_{p^{e}N} containing {±1}, let X_{G} be the smooth projective model of the quotient X(p^{e}; N)/G. The G_{p^{e}N} form an inverse system with inverse limit

S := ℤ_{p}^{×} × ∏_{prime ℓ ≠ p} SL_{2}(ℤ_{ℓ}).

The inverse image of G under S ↠ G_{p^{e}N} is an open subgroup of the profinite group S, and every open subgroup of S containing {±1} arises this way for some p^{e}N. Thus we may define X_{G} for any open subgroup G of S containing {±1}.

It seems likely that there is a constant c > 0 independent of p and G such that γ_{\overline{𝔽}_{p}}(X_{G}) ≥ c(S : G). We are unable to prove such a linear lower bound, even for fixed p, but we can show that the gonality goes to infinity for fixed p. Here is our main theorem:

Theorem 1.3. Fix a prime p. Let G_{1}, G_{2}, … be a sequence of distinct open subgroups of S containing \{±1\}. Then γ_{\overline{𝔽}_{p}}(X_{G_{i}}) → ∞ as i → ∞.

1.3. Outline of proof of main theorem. Many of the ideas used in the proof of Theorem 1.3 are due to earlier authors, though we consider a broader class of modular curves than had been treated earlier. Section 2 proves Theorem 2.5, an inequality in the direction opposite to Proposition A.1(i): the ideas used here and their application to the classical modular curves X_{0}(N) can be found in [HS91], [NS96], and [Bak99, Chapter 3]. Theorem 2.5 reduces the problem to finding lower bounds on gonality over finite fields, and these can be obtained by counting in the spirit of [Ogg74], which among other things determined the N for which X_{0}(N) is hyperelliptic. In Section 3 we find that, as in [Ogg74], modular curves of level prime to p have too many supersingular points over 𝔽_{p^{2}} to have small gonality. In Section 4 we cite results of Schweizer [Sch05], who obtained lower bounds on the \overline{𝔽}_{p}-gonality of Igusa curves directly from the geometry of the curves, instead of first getting lower bounds on 𝔽_{p}-gonality by counting 𝔽_{p}-points. Section 5 uses Goursat’s lemma to study the subgroups of S, so that the prime-to-p and p-power cases can be combined to prove the general case of Theorem 1.3 in Section 6.

1.4. Application to the image of Galois. One application of results like Theorem 1.3, noted already by many other authors, is to the function field analogue of the strong uniform boundedness theorem for elliptic curves. By the work of Mazur, Kamienny, and Merel [Mer96], for every d ∈ ℤ_{≥1}, there exists a constant N_{d} such that for any number field K with [K : ℚ] ≤ d and for any elliptic curve E over K, the torsion subgroup E(K)_{tors} of the finitely generated abelian group E(K) satisfies #E(K)_{tors} ≤ N_{d}. In the function field case, we can prove a stronger result, one which bounds the index of the image of Galois acting on torsion. If E is an ordinary elliptic curve over a field K of characteristic p ≥ 0, and K^{s} is a separable closure of K, there exists a homomorphism

ρ_{E}: Gal(K^{s}/K) → ℤ_{p}^{×} × ∏_{ℓ≠p} GL_{2}(ℤ_{ℓ})

describing the Galois action on \underleftarrow{lim} ker (V_{p^{e}}: E^{(p^{e})} → E) and the ℓ-adic Tate modules of E. (Of course, if char K = 0, there is no ℤ_{p}^{×} factor.)

Theorem 1.4. Given p ≥ 0 and d ∈ ℤ_{≥1}, there exists a constant N_{p,d} such that for any field k of characteristic p, any field K of degree ≤ d over k(t), and any elliptic curve E over K with j(E) not algebraic over k, the index (S : ρ_{E}(Gal(K^{s}/K)) ∩ S) is at most N_{p,d}.

Remark 1.5. Cojocaru and Hall [CH05, Theorem 1.1] give an explicit upper bound on the set of primes ℓ ≠ char k for which the image of ρ_{E} in GL_{2}(ℤ/ℓℤ) does not contain SL_{2}(ℤ/ℓℤ); their bound depends on the genus and not only on the gonality of the function field K.

Remark 1.6. In Theorem 1.4, we cannot hope to bound the index of ρ_{E}(Gal(K^{s}/K)) in ℤ_{p}^{×} × ∏_{ℓ≠p} GL_{2}(ℤ_{ℓ}) (i.e., with GL_{2} instead of the SL_{2} in the definition of S), since the determinant of the image in GL_{2}(ℤ_{ℓ}) gives the action of Gal(K^{s}/K) on roots of unity, and this is trivial if k is algebraically closed, for example.

Theorem 1.4 will be deduced from Theorem 1.3 in Section 7.

2. Change in gonality under extension of the ground field

In this section we give an exposition of the “tower theorem” of Nguyen and Saito [NS96, Theorem 2.1], and its implication for relating gonalities of a single curve over different fields. We will reprove it as our Proposition 2.4, since [NS96] remains unpublished after 10 years, and since we can simplify the proof slightly. Throughout this section, k is a perfect field.

Proposition 2.1 (Castelnuovo-Severi inequality). Let F, F_{1}, F_{2} be function fields of curves over k, of genera g, g_{1}, g_{2}, respectively. Suppose that F_{i} ⊆ F for i = 1, 2 and the compositum of F_{1} and F_{2} in F equals F. Let d_{i} = [F : F_{i}] for i = 1, 2. Then

g ≤ d_{1}g_{1} + d_{2}g_{2} + (d_{1} − 1)(d_{2} − 1).

Proof. See [Sti93, III.10.3]. □

Let X be a curve over k. A subfield F of k(X) will be called d-controlled if there exists e ∈ ℤ_{>0} such that [k(X) : F] = d/e and the genus of F is ≤ (e − 1)^{2}.

Lemma 2.2. If F is d-controlled, and f ∈ k(X) is a rational function of degree d, then F (f) is d-controlled.

Proof. View F (f) as the compositum of F and k(f) in k(X). Let a = [F (f) : F]. Then [k(X) : F (f)] = d/(ae), so [F (f) : k(f)] = ae. By Proposition 2.1, the genus of F (f) is at most

a(e − 1)^{2} + 0 + (a − 1)(ae − 1) = (ae − 1)^{2} − ae(a − 1)(e − 1) ≤ (ae − 1)^{2},

since a, e ≥ 1. □

Corollary 2.3. A subfield of k(X) generated over k by one or more elements of degree d is d-controlled.

Proof. Induction on the number of elements: the case of one element is trivial (e = 1), and Lemma 2.2 gives the inductive step. □

Proposition 2.4 (Tower theorem). Let X be a curve over a perfect field k. Let L ⊇ k be an algebraic field extension. Let d = γ_{L}(X). Then k(X) has a d-controlled subfield.

Proof. Enlarging L cannot increase γ_{L}(X), so we may assume L/k is Galois. Choose f ∈ L(X) of degree d. Let F_{L} be the subfield generated over L by the Gal(L/k)-conjugates of f. By Corollary 2.3, F_{L} is d-controlled as a subfield of L(X). The action of Gal(L/k) on L(X) preserves F_{L}, and the invariant subfield F_{k} := F_{L}^{Gal(L/k)} satisfies [k(X) : F_{k}] = [L(X) : F_{L}] and has the same genus as F_{L}. Thus F_{k} is a d-controlled subfield of k(X). □

Theorem 2.5. Let X be a curve over a perfect field k. Let L ⊇ k be an algebraic field extension. Let d = γ_{L}(X). Assume that X(k) ≠ ∅.

(i) If d ≤ 2, then γ_{k}(X) = d.

(ii) If d > 2, then γ_{k}(X) ≤ (d − 1)^{2}.

(iii) In any case, γ_{L}(X) ≥ \sqrt{γ_{k}(X)}.

Proof.

(i) If d = 1, then X ≃ ℙ_{k}^{1}, so γ_{k}(X) = 1. If d = 2, then X is elliptic or hyperelliptic; if elliptic, then γ_{k}(X) = 2; if hyperelliptic then the canonical map is a degree-2 map to a genus-0 curve Z over k, and Z(k) ≠ ∅ so Z ≃ ℙ_{k}^{1}, so γ_{k}(X) = 2.

(ii) Now suppose d > 2. By Proposition 2.4 there exists e ∈ ℤ_{>0} and a rational map π: X −−→ Y of curves over k such that deg π = d/e and the genus g of Y satisfies g ≤ (e − 1)^{2}. We have Y (k) ≠ ∅. If g = 0, then Y ≃ ℙ^{1}, so

γ_{k}(X) ≤ d/e ≤ d < (d − 1)^{2}.

If g = 1, then e ≥ 2 and γ_{k}(Y) = 2, so

γ_{k}(X) ≤ (d/e)γ_{k}(Y) ≤ (d/2)2 = d < (d − 1)^{2}.

If g ≥ 2, then γ_{k}(Y) ≤ g by Proposition A.1(iv), so

γ_{k}(X) ≤ \frac{d}{e}γ_{k}(Y) ≤ \frac{d}{e}(e − 1)^{2}.

For e ∈ [1, d], the function \frac{d}{e}(e − 1)^{2} is maximized at e = d, and the value there is (d − 1)^{2}.

(iii) This follows directly from the first two parts.

□

Remark 2.6. The hypothesis X(k) ≠ ∅ is necessary: Genus-1 curves over ℚ have \overline{ℚ}-gonality 2, but their ℚ-gonality can be arbitrarily large.

Remark 2.7. We do not know whether the (d − 1)^{2} in Theorem 2.5(ii) can be improved.

Remark 2.8. For N = 38, 44, 53, 61, the modular curve X_{0}(N) is of genus 4 and has ℚgonality 4 and \overline{ℚ}-gonality 3 [HS99, p. 136]. In particular, Theorem 2.5(ii) is best possible for d = 3.

3. Level prime to p

Suppose p ∤ N. We begin by defining a twisted form X(N)′ over 𝔽_{p^{2}} of X(N). Let M be (ℤ/Nℤ)^{2} made into a Gal(\overline{𝔽}_{p}/𝔽_{p^{2}})-module by letting the p^{2}-power Frobenius automorphism act as multiplication by −p. There exists an isomorphism of Galois modules ι: ⋀^{2} M → μ_{N}; fix one. Let X(N)′ be the smooth projective model of the affine curve over 𝔽_{p^{2}} parameterizing pairs (E, ϕ) where E is an elliptic curve and ϕ is an isomorphism E[N] → M under which the Weil pairing corresponds to ι. Over \overline{𝔽}_{p}, X(N)′ becomes isomorphic to X(N). The automorphisms of M as an abelian group automatically commute with the Galois action, so they induce automorphisms of X(N)′ defined over 𝔽_{p^{2}}. Thus we get SL_{2}(ℤ/Nℤ)/{±1} ≤ Aut X(N)′. Moreover, it follows from [BGJGP05, Lemma 3.21] that all the points of X(N)′ corresponding to supersingular elliptic curves are defined over 𝔽_{p^{2}}.

Proposition 3.1. Let p, N, and X(N)′ be as above. Let G be a subgroup of SL_{2}(ℤ/Nℤ)/{±1} of index D. Let X be the curve X(N)′/G. Then the 𝔽_{p^{2}}-gonality γ of X satisfies

γ ≥ \frac{p − 1}{12(p^{2} + 1)}D.

Proof. This resembles the proof of [BGJGP05, Lemma 3.22]. By [BGJGP05, Lemma 3.20], the number of supersingular points on X is ≥ (p − 1)D/12, and these are images of supersingular points on X(N)′ so they are defined over 𝔽_{p^{2}}; thus #X(𝔽_{p^{2}}) ≥ (p − 1)D/12. On the other hand, #X(𝔽_{p^{2}}) ≤ γ#ℙ^{1}(𝔽_{p^{2}}) = γ(p^{2} + 1). Combine the two previous sentences. □

Remark 3.2. Let g be the genus of X. One could also combine Proposition 3.1 with the bound g − 1 < D/12 of [Shi94, Proposition 1.40] to give a lower bound for γ in terms of g instead of D.

We now consider the \overline{𝔽}_{p}-gonality of all modular curves of level prime to p.

Corollary 3.3. Fix p. Define Φ_{p}(D) := \sqrt{\frac{p−1}{12(p^{2}+1)}D}. If G is the inverse image under S ↠ ∏_{ℓ≠p} SL_{2}(ℤ_{ℓ}) of an open subgroup of index D in ∏_{ℓ≠p} SL_{2}(ℤ_{ℓ}) containing {±1}, then γ_{\overline{𝔽}_{p}}(X_{G}) ≥ Φ_{p}(D).

Proof. Combine Proposition 3.1 and Theorem 2.5(iii). □

Remark 3.4. One could also consider Atkin-Lehner quotients of X_{0}(N) for N prime to p. These are generally not of the form X_{G}. Nevertheless, gonality bounds tending to infinity for fixed p can be obtained: first apply Proposition 3.1 to get a lower bound on γ_{𝔽_{p^{2}}}(X_{0}(N)), next use Proposition A.1(vi) to get a lower bound on the 𝔽_{p^{2}}-gonality of any quotient of X_{0}(N), and finally apply Theorem 2.5. This works since the size of the Atkin-Lehner group (some power of 2) is asymptotically small compared to the index of the congruence subgroup Γ_{0}(N) in PSL_{2}(ℤ).

4. Level a power of p

The necessary lower bound on the gonality of Ig(p^{e}) has been proved already by Schweizer, in a strong form:

Theorem 4.1 ([Sch05, Lemma 1.5(d,e)]).

(i) If p ≥ 7, then \frac{p+13}{24} ≤ γ_{\overline{𝔽}_{p}}(Ig(p)) ≤ \frac{p−1}{6}.

(ii) If e > 1 and p^{e} ∉ \{25, 9, 8, 4\}, then γ_{\overline{𝔽}_{p}}(Ig(p^{e})) = pγ_{\overline{𝔽}_{p}}(Ig(p^{e−1})).

In fact, [Sch05] proves many more results. The above are more than we need to deduce the following.

Corollary 4.2. Let G be the inverse image under S ↠ ℤ_{p}^{×} of an open subgroup of index D in ℤ_{p}^{×} containing \{±1\}. Then γ_{\overline{𝔽}_{p}}(X_{G}) > D/12.

Proof. First suppose that X_{G} = Ig(p^{e}) for some prime power p^{e} > 2. Then D = p^{e−1}(p−1)/2. For p^{e} ≤ 25, we have D < 12, and γ_{\overline{𝔽}_{p}}(X_{G}) ≥ 1 > D/12 trivially. For p > 25, Theorem 4.1(i) gives

γ_{\overline{𝔽}_{p}}(Ig(p)) ≥ \frac{p + 13}{24} > \frac{p − 1}{24} = \frac{D}{12}.

For p^{e} > 25 with e > 1, we use induction on e, with Theorem 4.1(ii) giving the inductive step.

Any other X_{G} in Corollary 4.2 is a quotient of Ig(p^{e}) for some p^{e} > 2, and the inequality for X_{G} follows from the inequality for Ig(p^{e}), by Proposition A.1(vi). □

Remark 4.3. An alternative approach to lower bounds on the gonality of Igusa curves is to show that they have many points over certain finite fields, to deduce that the gonality over these finite fields is large, and then to apply Theorem 2.5. One can no longer use supersingular points, however, since these are totally ramified in Ig(p^{e}) →^{j} ℙ^{1}, and hence their number does not grow with e. Instead we could use ordinary points: it follows from [Pac96, Corollary 2.13] and Hurwitz class number estimates that # Ig(q)(𝔽_{q}) ≥ q^{3/2−o(1)} as q → ∞. Or one could use cusps, as in the proof of [Sch04, Theorem 6.1], to get lower bounds on # Ig(q)(𝔽_{p}), since the cusps split completely in Ig(p^{e}) →^{j} ℙ^{1} [KM85, Corollary 12.7.2]. But the lower bounds on gonality obtained by these methods are weaker than the ones we took from [Sch05].

5. Group theory

Here we study the open subgroups of S. Let S_{p} = ℤ_{p}^{×} and S_{≠p} = ∏_{ℓ≠p} SL_{2}(ℤ_{ℓ}), so S = S_{p} × S_{≠p}. Given an open subgroup G ≤ S, let G_{p} and G_{≠p} be the images of G in S_{p} and S_{≠p}, respectively.

Lemma 5.1. Let B ∈ ℤ_{>0}. Any open subgroup H of S_{≠p} of index ≤ B contains

∏_{ℓ≤B!, ℓ≠p}{1} × ∏_{ℓ>B!, ℓ≠p} SL_{2}(ℤ_{ℓ}).

Proof. For each ℓ ≠ p, identify SL_{2}(ℤ_{ℓ}) with a subgroup of S_{≠p} in the obvious way. It suffices to show that H contains SL_{2}(ℤ_{ℓ}) for each ℓ > B! with ℓ ≠ p.

The kernel of the action of S_{≠p} on the coset space S_{≠p}/H is a normal open subgroup N ⊴ S_{≠p} contained in H. Let n := (S_{≠p} : N), so n ≤ B!. Now ℓ > B! ≥ n, so 1/n ∈ ℤ_{ℓ}, and

(\begin{matrix}1 & 1 \\ 0 & 1 \end{matrix}) = (\begin{matrix}1 & 1/n \\ 0 & 1 \end{matrix})^{n} ∈ N,

where the matrices belong to SL_{2}(ℤ_{ℓ}) ≤ S_{≠p}. Similarly (\begin{matrix}1 & 0 \\ 1 & 1 \end{matrix}) ∈ N. But these two matrices

generate the dense subgroup SL_{2}(ℤ) of SL_{2}(ℤ_{ℓ}), so SL_{2}(ℤ_{ℓ}) ≤ N ≤ H. □

Lemma 5.2. For each B > 0, there are at most finitely many open subgroups G of S such that (S_{p} : G_{p}) < B and (S_{≠p} : G_{≠p}) < B.

Proof. Fix B. By Lemma 5.1, it suffices to consider instead the situation in which S_{≠p} is replaced by S_{L} := ∏_{ℓ∈L} SL_{2}(ℤ_{ℓ}) for a finite set L of primes ≠ p: i.e., G is now an open subgroup of S_{p} × S_{L}, G_{L} is the image of G in S_{L}, and we are given (S_{p} : G_{p}) < B and (S_{L} : G_{L}) < B.

Since S_{p} and S_{≠p} are topologically finitely generated, there are finitely many possibilities for G_{p} and G_{L}. Goursat’s Lemma [Lan02, p. 75] states that each possible G is the inverse image under

G_{p} × G_{L} ↠ \frac{G_{p}}{H_{p}} × \frac{G_{L}}{H_{L}}

of the graph of an isomorphism

\frac{G_{p}}{H_{p}} → \frac{G_{L}}{H_{L}}

for some normal open subgroups H_{p} ⊴ G_{p} and H_{L} ⊴ G_{L}. By the finite generation again, it suffices to bound (G_{p} : H_{p}) = (G_{L} : H_{L}). It is bounded by the supernatural number gcd(#S_{p}, #S_{L}), which is finite, since S_{p} has a pro-p open subgroup, while S_{L} has an open subgroup of order prime to p. (See [Ser02, I.§1.3] for the notion of supernatural number.) □

6. The general case of Theorem 1.3

Proof of Theorem 1.3. Let m > 0. We will show that γ_{\overline{𝔽}_{q}}(X_{G_{i}}) > m for all but finitely many i. Let Φ_{p} be as in Corollary 3.3. Choose B_{0} such that min{Φ_{p}(B), B/12} > m for all B ≥ B_{0}. By Lemma 5.2, all but finitely many G_{i} in our sequence have either (S_{p} : (G_{i})_{p}) ≥ B or (S_{≠p} : (G_{i})_{≠p}) ≥ B. For each such i, X_{G_{i}} dominates an X_{G} with G as in Corollary 3.3 or Corollary 4.2; then γ_{\overline{𝔽}_{p}}(X_{G}) exceeds either Φ_{p}(B) or B/12. By Proposition A.1(vii),

γ_{\overline{𝔽}_{p}}(X_{G_{i}}) ≥ γ_{\overline{𝔽}_{p}}(X_{G}) ≥ min\{Φ_{p}(B), B/12\} > m.
□

7. Image of Galois

Proof of Theorem 1.4. We may assume that k is algebraically closed. By Theorem 1.3, there are only finitely many subgroups G ≤ S containing \{±1\} such that γ_{\overline{𝔽}_{p}}(X_{G}) ≤ d. Choose N_{p,d} such that N_{p,d} ≥ 2(S : G) for every such G.

Let K be a field of degree ≤ d over k(t), and let E be an elliptic curve over K with j(E) not algebraic over k (i.e., not in k). Then E is ordinary. Write K = k(C), where C is a curve over k with γ_{k}(C) ≤ d. Define H := ρ_{E}(Gal(K^{s}/K)). Since k is algebraically closed, H ⊆ S. We want (S : H) ≤ N_{p,d}.

Suppose not. If (S : H) is infinite, then since S/H is a profinite group, we can find a group H′ with H ≤ H′ ≤ S and N_{p,d} < (S : H′) < ∞. If (S : H) is finite, let H′ = H. In either case, let H′′ be the group generated by H′ and −1, so N_{p,d}/2 < (S : H′′) < ∞.

By definition of N_{p,d}, the group H′′ does not equal any of the groups G, so γ_{\overline{𝔽}_{p}}(X_{H′′}) > d. Equivalently, by Proposition A.1(ii), γ_{k}(X_{H′′}) > d.

The curve X_{H′′} is defined as a quotient of some X(p^{e}; N). Choosing level structure for E over K^{s} gives a point in X(p^{e}; N)(K^{s}), and the action of Gal(K^{s}/K) moves this point within the H′′-orbit, since H ⊆ H′′, so the image point in X_{H′′}(K^{s}) is K-rational. This point in X_{H′′}(K) may be viewed as a rational map C −−→ X_{H′′}, and this map is non-constant since the composition C −−→ X_{H′′} →^{j} X(1) ≃ ℙ^{1} corresponds to j(E) ∈ K − k. Proposition A.1(vii) implies γ_{k}(X_{H′′}) ≤ γ_{k}(C) ≤ d, contradicting the previous paragraph. □

Appendix A. General facts about gonality

Proposition A.1. Let X be a curve of genus g over a field k.

(i) If L is a field extension of k, then γ_{L}(X) ≤ γ_{k}(X). (ii) If k is algebraically closed, and L is a field extension of k, then γ_{L}(X) = γ_{k}(X). (iii) If g > 1, then γ_{k}(X) ≤ 2g −2. For each g > 1, there exist k and X for which equality

holds. (iv) If X(k) ≠ ∅, then γ_{k}(X) ≤ g + 1. If X(k) ≠ ∅ and g ≥ 2, then γ_{k}(X) ≤ g (and

again, equality is possible for each g).

(v) If k is algebraically closed, then γ_{k}(X) ≤ ⌊\frac{g+3}{2}⌋. Equality holds for a general curve of genus g over k.

(vi) If π: X −−→ Y is a dominant rational map of curves over k, then γ_{k}(X) ≤ (deg π)γ_{k}(Y). (vii) If π: X −−→ Y is a dominant rational map of curves over k, then γ_{k}(Y) ≤ γ_{k}(X).

Proof.

(i) Trivial. (ii) Given a map over L, standard specialization arguments give a map over k of the

same degree. (iii) The canonical linear system |K| has dimension g − 1 ≥ 1 and degree 2g − 2. So we

may use a rational function whose divisor is the difference of two different canonical divisors. Equality holds for the general curve over the function field of the moduli space M_{g} of genus-g curves in characteristic 0, since its only line sheaves are the powers of the canonical sheaf: this was the Franchetta conjecture, proved in [Har83] and strengthened in [Mes87].

(iv) Let P ∈ X(k). The Riemann-Roch theorem shows that dim |(g + 1)P | ≥ 1, and that dim |K − (g − 2)P | ≥ 1 if g − 2 ≥ 0. These linear systems have degree g + 1 and g, respectively.

Now we prove that equality is possible, by considering the general curve X_{g,1} over the function field of the moduli space M_{g,1} of genus-g curves with one marked point in characteristic 0. We may assume g ≥ 3. The Picard group of X_{g,1} is generated by the canonical sheaf K and the class of the marked point P: this extension of the Franchetta conjecture can be deduced from the description of the Picard group of the moduli stack of genus-g curves with two marked points, in the same way that the original Franchetta conjecture is deduced in [AC87, §4]. It remains to show that dim |aK + bP | < 1 whenever a, b ∈ ℤ are such that deg(aK + bP) < g. By adding multiples of P, it suffices to consider the case deg(aK + bP) = g − 1. By Riemann-Roch symmetry, we may assume a ≥ 1. If a = 1, we have dim |K − (g − 1)P | = 0 since the general point P is not a Weierstrass point. If a > 1, |aK −(2a−1)(g −1)P | is empty by [Nee84, Theorem 4.2].

(v) These are consequences of Brill-Noether theory: see (1.1) and (1.5) of [ACGH85, Chapter V] for an exposition. The first statement is proved in arbitrary characteristic in [KL72,KL74]. The second statement is proved in characteristic 0 in [Far66,Mar67, Mar68], and can be deduced in characteristic p from the unramified case of [Oss05, Theorem 1.2], for instance.

(vi) Trivial. (vii) (The ideas in the following argument go back at least to [New72, Theorem VII.2].)

Choose f ∈ k(X) of degree d := γ_{k}(X). Let r = deg π = [k(X) : k(Y)]. Let P (T) ∈ k(Y)[T] be the characteristic polynomial of f viewed as an element of the field extension k(X) of k(Y). For some finite normal extension M of k(Y), we may write P (T) = ∏_{i=1}^{r}(T −f_{i}) for some f_{i} ∈ M. As a function in M, f has degree [M : k(X)]d. The same is true of each f_{i}, since they are all in the same Aut(M/k(Y))-orbit. The polar divisor of a coefficient of P viewed in M is at most the sum of the polar divisors of the f_{i}, so each coefficient has degree at most r[M : k(X)]d = [M : k(Y)]d as a function in M, and hence degree at most d as a function in k(Y). Since f is non-constant, at least one of these coefficients is non-constant. Thus γ_{k}(Y) ≤ d.

□

Acknowledgements

Most of all I thank Andreas Schweizer for several discussions, for suggesting references, and for pointing out that existing versions of the Castelnuovo-Severi inequality seem to require a perfect field of constants. I also thank Matt Baker, from whom I first learned the idea in [HS91] that the Castelnuovo-Severi inequality could be used to bound change in gonality under algebraic extensions. I thank Brian Osserman for pointing out that the lower bound on the geometric gonality of the general curve of genus g in characteristic p could be deduced from [Oss05]. I thank Joe Harris for suggesting the reference [AC87] and Doug Ulmer for a remark about cusps of Igusa curves, Finally I thank Alina Cojocaru, Chris Hall, Dinesh Thakur, and Doug Ulmer for asking questions that inspired this paper.

References

[Abr96] Dan Abramovich, A linear lower bound on the gonality of modular curves, Internat. Math. Res. Notices (1996), no. 20, 1005–1011.MR1422373 (98b:11063) ↑1.1, 1.2

[AC87] Enrico Arbarello and Maurizio Cornalba, The Picard groups of the moduli spaces of curves, Topology 26 (1987), no. 2, 153–171. MR 895568 (88e:14032) ↑iv, A

[ACGH85] E. Arbarello, M. Cornalba, P. A. Griffiths, and J. Harris, Geometry of algebraic curves. Vol. I, Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], vol. 267, Springer-Verlag, New York, 1985.MR770932 (86h:14019) ↑v

[Bak99] Matthew H. Baker, Torsion points on modular curves, 1999. Ph.D. thesis, University of California at Berkeley. ↑1.3

[BGJGP05] Matthew H. Baker, Enrique González-Jiménez, Josep González, and Bjorn Poonen, Finiteness results for modular curves of genus at least 2, Amer. J. Math. 127 (2005), 1325–1387.

   arXiv:math.NT/0211394. ↑1.2, 3, 3

[CH05] Alina Carmen Cojocaru and Chris Hall, Uniform results for Serre’s theorem for elliptic curves, Int. Math. Res. Not. (2005), no. 50, 3065–3080. MR 2189500 ↑1.5 [Far66] Hershel M. Farkas, Special divisors and analytic subloci of Teichmueller space, Amer. J. Math.

88 (1966), 881–901.MR0213546 (35 #4406) ↑v

[Har83] John Harer, The second homology group of the mapping class group of an orientable surface, Invent. Math. 72 (1983), no. 2, 221–239.MR700769 (84g:57006) ↑iii

[HS91] Joe Harris and Joe Silverman, Bielliptic curves and symmetric products, Proc. Amer. Math. Soc. 112 (1991), no. 2, 347–356.MR1055774 (91i:11067) ↑1.3, A

[HS99] Yuji Hasegawa and Mahoro Shimura, Trigonal modular curves, Acta Arith. 88 (1999), no. 2, 129–140.MR1700245 (2000d:11080) ↑2.8

[KM85] Nicholas M. Katz and Barry Mazur, Arithmetic moduli of elliptic curves, Annals of Mathematics Studies, vol. 108, Princeton University Press, Princeton, NJ, 1985.MR772569 (86i:11024) ↑4.3

[KL72] Steven L. Kleiman and Dan Laksov, On the existence of special divisors, Amer. J. Math. 94 (1972), 431–436.MR0323792 (48 #2148) ↑v

[KL74] , Another proof of the existence of special divisors, Acta Math. 132 (1974), 163–176.MR0357398 (50 #9866) ↑v

[Lan02] Serge Lang, Algebra, 3rd ed., Graduate Texts in Mathematics, vol. 211, Springer-Verlag, New York, 2002.MR1878556 (2003e:00003) ↑5

[Mar67] Henrik H. Martens, On the varieties of special divisors on a curve, J. Reine Angew. Math. 227 (1967), 111–120.MR0215847 (35 #6682) ↑v

[Mar68] , Varieties of special divisors on a curve. II, J. Reine Angew. Math. 233 (1968), 89–100.MR0241420 (39 #2760) ↑v

[Mer96] Loïc Merel, Bornes pour la torsion des courbes elliptiques sur les corps de nombres, Invent. Math. 124 (1996), no. 1-3, 437–449 (French).MR1369424 (96i:11057) ↑1.4

[Mes87] Nicole Mestrano, Conjecture de Franchetta forte, Invent. Math. 87 (1987), no. 2, 365–376 (French).MR870734 (88c:14039) ↑iii

[Nee84] A. Neeman, Weierstrass points in characteristic p, Invent. Math. 75 (1984), no. 2, 359–376. MR 732551 (85h:14009) ↑iv

[New72] Morris Newman, Conjugacy, genus, and class numbers, Math. Ann. 196 (1972), 198–217.MR0311573 (47 #135) ↑vii

[NS96] Khac Viet Nguyen and Masa-Hiko Saito, d-gonality of modular curves and bounding torsions (March 29, 1996). Preprint, arXiv:alg-geom/9603024. ↑1.3, 2

[Ogg74] Andrew P. Ogg, Hyperelliptic modular curves, Bull. Soc. Math. France 102 (1974), 449–462.MR0364259 (51 #514) ↑1.3

[Oss05] Brian Osserman, Deformations of covers, Brill-Noether theory, and wild ramification, Math. Res. Lett. 12 (2005), no. 4, 483–491.MR2155226 ↑v, A

[Pac96] Amílcar Pacheco, Rational points on Igusa curves and L-functions of symmetric representations, J. Number Theory 58 (1996), no. 2, 343–360.MR1393620 (97e:11078) ↑4.3

[Sch04] Andreas Schweizer, Torsion of Drinfeld modules and gonality, Forum Math. 16 (2004), no. 6, 925–941.MR2096477 (2005f:11116) ↑4.3

[Sch05] , On the p^{e}-torsion of elliptic curves and elliptic surfaces in characteristic p, Trans. Amer. Math. Soc. 357 (2005), no. 3, 1047–1059 (electronic).MR2110432 (2005k:11111) ↑1.3, 4.1, 4, 4.3

[Ser02] Jean-Pierre Serre, Galois cohomology, Corrected reprint of the 1997 English edition, Springer Monographs in Mathematics, Springer-Verlag, Berlin, 2002. Translated from the French by Patrick Ion and revised by the author.MR1867431 (2002i:12004) ↑5

[Shi94] Goro Shimura, Introduction to the arithmetic theory of automorphic functions, Publications of the Mathematical Society of Japan, vol. 11, Princeton University Press, Princeton, NJ, 1994. Reprint of the 1971 original; Kanô Memorial Lectures, 1.MR1291394 (95e:11048) ↑1.2, 3.2

[Sti93] Henning Stichtenoth, Algebraic function fields and codes, Universitext, Springer-Verlag, Berlin, 1993.MR1251961 (94k:14016) ↑2

Department of Mathematics, University of California, Berkeley, CA 94720-3840, USA

E-mail address: poonen@math.berkeley.edu
URL: http://math.berkeley.edu/~poonen